# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v1

@@ Appendix B Proofs
It immediately follows from Theorem B.4 that $[\mathbb{P}^{0}]^{N}(\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N}))\geq 1-\eta_{1}$ holds. If we take $\widehat{\varepsilon}_{1}>0$ as
@@
The second statement immediately follows under the assumptions of Theorem B.8Inter-ARO overestimates the true loss analogously as Theorem B.5 with an identical proof.
Appendix C Exponential cone representation of DR-ARO
@@
Following the conic modeling guidelines of [41], for new variables $v^{+}_{i},w^{+}_{i}\in\mathbb{R}$, the first constraint can be written as
@@
Appendix D Further details for numerical experiments
All experiments are implemented in Julia [11] (MIT license) and executed on Intel Xeon 2.66GHz processors with 8GB memory in single-core mode. We use MOSEK 10.1 [42] to solve all exponential conic programs through JuMP [23]. The UCI datasets [22, 39] we use (see Table 4) are subject to CC BY 4.0 license. MNIST is subject to CC BY-SA 3.0 and EMNIST to CC0 1.0 license.
D.1 UCI experiments
@@
Although we reported the first 5 datasets in the main paper, we experiment on 10 UCI datasets [22, 39] (cf. Table 4). We use Python3 for preprocessing these datasets. Classification problems with more than two classes are converted to binary classification problems (most frequent class/others). For all datasets, numerical features are standardized, the ordinal categorical features are left as they are, and the nominal categorical features are processed via one-hot encoding. As mentioned in the main paper, we obtain auxiliary (synthetic) datasets via SDV, which is also implemented in Python 3.
@@
Finally, we demonstrate that our theory, especially DRO+ARO+Aux, contributes to the DRO literature even without adversarial attacks. In this case of $\alpha=0$, ERM and ARO would be equivalent, and DRO+ARO would reduce to the traditional DR LR model [60]. ARO+Aux would be interpreted as revising the empirical distribution of ERM to a mixture (mixture weight cross-validated) of the empirical and auxiliary distributions. DRO+ARO+Aux, on the other hand, can be interpreted as DRO over a carefully reduced ambiguity set (intersection of the empirical and auxiliary Wasserstein balls). The results are in Table 6. Analogous results follow as before (that is, DRO+ARO+Aux is the ‘winning’ approach, DRO+ARO and ARO+Aux alternate for the ‘second’ approach), with the exception of the dataset contraceptive, where ARO+Aux outperforms others.
@@
D.2 MNIST/EMNIST experiments

### v2

@@ 10.10 Proof of Theorem 3
It immediately follows from Theorem 4 that $[\mathbb{P}^{0}]^{N}(\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N}))\geq 1-\eta_{1}$ holds. If we take $\widehat{\varepsilon}_{1}>0$ as
@@
The second statement immediately follows under the assumptions of Theorem 8Inter-ARO overestimates the true loss analogously as Theorem 5 with an identical proof.
11 EXPONENTIAL CONIC REFORMULATION OF DR-ARO
@@
Following the conic modeling guidelines of MOSEK ApS, (2023), for new variables $v^{+}_{i},w^{+}_{i}\in\mathbb{R}$, the first constraint can be written as
@@
12 FURTHER DETAILS ON NUMERICAL EXPERIMENTS
12.1 UCI Experiments
@@
We experiment on 10 UCI datasets (Kelly et al.,, 2023) (cf. Table 3). We use Python 3 for preprocessing these datasets. Classification problems with more than two classes are converted to binary classification problems (most frequent class/others). For all datasets, numerical features are standardized, the ordinal categorical features are left as they are, and the nominal categorical features are processed via one-hot encoding. As mentioned in the main paper, we obtain auxiliary (synthetic) datasets via SDV, which is also implemented in Python 3.
@@
Finally, we demonstrate that our theory, especially DRO+ARO+Aux, contributes to the DRO literature even without adversarial attacks. In this case of $\alpha=0$, ERM and ARO would be equivalent, and DRO+ARO would reduce to the traditional DR LR model (Shafieezadeh-Abadeh et al.,, 2015). ARO+Aux would be interpreted as revising the empirical distribution of ERM to a mixture (mixture weight cross-validated) of the empirical and auxiliary distributions. DRO+ARO+Aux, on the other hand, can be interpreted as DRO over a carefully reduced ambiguity set (intersection of the empirical and auxiliary Wasserstein balls). The results are in Table 5. Analogous results follow as before (that is, DRO+ARO+Aux is the ‘winning’ approach, DRO+ARO and ARO+Aux alternate for the ‘second’ approach), with the exception of the dataset contraceptive, where ARO+Aux outperforms others.
@@
12.2 MNIST/EMNIST Experiments